# ROOFS: RObust biOmarker Feature Selection

### v1

@@ Abstract
Feature selection (FS) is essential for biomarker discovery and in the analysis of biomedical datasets. However, challenges such as high-dimensional feature space, low sample size, multicollinearity, and missing values make FS non-trivial. Moreover, FS performances vary across datasets and predictive tasks. We propose roofs, a Python package available at https://gitlab.inria.fr/compo/roofs, designed to help researchers in the choice of FS method adapted to their problem. roofs benchmarks multiple FS methods on the user’s data and generates reports that summarize a comprehensive set of evaluation metrics, including downstream predictive performance estimated using optimism correction, stability, reliability of individual features, and true positive and false positive rates assessed on semi-synthetic data with a simulated outcome. We demonstrate the utility of roofs on data from the PIONeeR clinical trial, aimed at identifying predictors of resistance to anti-PD-(L)1 immunotherapy in lung cancer. The PIONeeR dataset contained 374 multi-source blood and tumor biomarkers from 435 patients. A reduced subset of 214 features was obtained through iterative variance inflation factor pre-filtering. Of the 34 FS methods gathered in roofs, we evaluated 23 in combination with 11 classifiers (253 models in total) and identified a filter based on the union of Benjamini-Hochberg false discovery rate-adjusted p-values from t-test and logistic regression as the optimal approach, outperforming other methods including the widely used LASSO. We conclude that comprehensive benchmarking with roofs has the potential to improve the robustness and reproducibility of FS discoveries and increase the translational value of clinical models.
@@ 1 Introduction
Feature selection (FS) refers to a class of methods aimed at reducing the feature space by selecting a subset of size $k<p$. In supervised learning, FS is used either for discovery (i.e., to identify the most informative variables), or for prediction, to develop models with reduced variance. A basic requirement for FS is that the reduced model should perform equivalently to or better than the full model, capturing the complexity of the original data. In the context of biomarker discovery and clinical predictive models, the selected signature, defined as a minimal set of variables with high predictive capacity, should meet additional requirements: (i) contain biologically or clinically interpretable features; (ii) be as minimal as possible; and (iii) demonstrate stability, i.e. the selected features should not be sensitive to small changes in the data [28]. FS stability is particularly critical for ensuring reproducibility of biomarker signatures. It is typically assessed by measuring the similarity between feature subsets obtained across resamples of the original data, e.g. via bootstrapping or subsampling. A method that is not able to reproduce similar feature subsets across resamples is unlikely to capture a reliable signature in the original dataset and may result in models that do not generalize well to external populations. Consequently, risk prediction models or biomarker panels built on unstable signatures are unlikely to be clinically useful. Importance of stability evaluation has been frequently acknowledged in the literature [28, 20, 22] but is yet often neglected in practice.
The complexity of biomedical data has several important implications for FS. First, the number of samples (here, number of patients, denoted $n$) is often smaller than the number of features (denoted $p$), leading to ill-posedness. Second, the often-encountered high degree of multicollinearity makes FS prone to instability because correlated features can be used interchangeably by a model. Third, high rates of missing values are common, which further complicates the discovery, increasing the rate of false positives and false negatives [41].
@@
A wide range of FS methods has been proposed in the literature (see [31] for an extensive review). Briefly, based on the underlying algorithmic approach, FS methods can be classified into 4 groups: filter, embedded, wrapper, and ensemble methods. Filters evaluate features according to a specific criteria and output either a complete ranking or a subset selected using a predefined threshold (e.g., univariable or multivariable statistical tests). Embedded methods perform FS simultaneously with model learning, as in regularization-based approaches or tree-based models with constrained depth. Wrapper methods iteratively search for feature subsets that maximize the predictive performance of a given model. Ensemble methods aim to improve the robustness of FS by aggregating outputs from different selection algorithms or from repeated runs of a single method on resampled data. Filter and embedded methods are typically more straightforward, computationally efficient, and interpretable compared to wrapper and ensemble techniques. At the same time, previous benchmarks [4, 13] have shown that, in terms of predictive performances, the optimal method is contingent upon the specific dataset, and no FS method consistently outperformed others across diverse predictive tasks. Furthermore, FS performance is sensitive to hyperparameter choice, such as the selected subset size, which could alter the relative ranking of methods within the same dataset [43].
@@
Given the numerous challenges in FS for biomedical data highlighted above, researchers may be hesitant about which FS approach to use when analyzing a new dataset. To address this gap, we propose a Python package called roofs available at https://gitlab.inria.fr/compo/roofs. roofs conducts comprehensive benchmarking of multiple pre-implemented FS algorithms on the user’s data based on three criteria: 1) method stability and robustness of individual features, 2) predictive performances, and 3) true positive and false positive rates (from semi-synthetic data with a simulated outcome). For point 2), we relied on the optimism-correction (OC) framework emphasized as a guideline for development of clinical predictive models by Collins et al. [12]. roofs provides automated reporting of the evaluation metrics listed above – a key feature of the package aimed to guide users in selecting the FS approach adapted to their dataset.
@@
We demonstrate the practical value of roofs by applying it to a complex dataset from the PIONeeR clinical study (NCT03833440, Precision Immuno-Oncology for Advanced Non-small Cell Lung Cancer Patients With PD-1 ICI Resistance) to predict the resistance to immunotherapy. This dataset is an interesting and challenging use case for FS because: 1) the data is rich and heterogeneous combining clinical features, routine blood tests and detailed immunoprofiling of blood and tumor samples measured by different clinical and research partners; 2) it contains clusters of highly correlated features, e.g. numerous variables describing related immune cell populations; and 3) $p$ (374) is roughly equivalent to $n$ (435). This is different from the three types of datasets popular in the literature: a) high-dimensional omics ($p\gg n$); b) data with a sufficiently high event-per-variable ratio, e.g. the number of patients in the smallest class is at least ten times the number of predictors $n_{min}\geq 10p$ (for classification problems); and c) big data (e.g., from electronic health records), typically modeled using deep artificial neural networks.
@@
A comprehensive survey of FS benchmarks across multiple domains published before 2020 is provided by Hopf and Reifenrath [23]. Table 1 summarizes more recent benchmarks on high-dimensional biomedical data. Most of these studies focused on a single data modality (genomic data). Only two studies included FS methods representing all algorithmic families, and stability was evaluated in only three of the eight benchmarks. The strengths of our study on PIONeeR are the following: 1) it addresses heterogeneous, mixed-type data; 2) it evaluates a broader and more diverse range of FS algorithms; 3) it reports all standard performance metrics, with an emphasis on stability; and 4) it extends FS evaluation to assess the balance between true and false discoveries.
To summarize our contribution:
1) In the Methods section, we introduce a Python framework for comprehensive evaluation of FS methods on user-provided data to help identify the most suitable FS approach for a given dataset and predictive task.
2) In the Results section, we demonstrate the applicability of this framework using a clinical dataset from lung cancer patients comprising diverse biomarkers, including demographic, clinical, and immunoprofiling data from both blood and tumor samples, to improve the prediction of resistance to immunotherapy.
@@ F E W Ens Predictive performance Stability Runtime
Our work Clinical + biological 1 23
@@ 2.1.1 PIONeeR dataset
The development of roofs was motivated by the following real-world problem: identifying robust biomarkers of resistance to immunotherapy in lung cancer. To this end, we relied on the PIONeeR clinical trial dataset, derived from a prospective, multicenter cohort of patients (17 centers in France) with advanced or recurrent non-small cell lung cancer (NSCLC). Patients received either frontline combination therapy with platinum-based chemotherapy plus anti-PD-(L)1 immune checkpoint inhibitors (ICIs) or second-/third-line monotherapy with anti-PD-(L)1 ICIs following progression on prior platinum-based chemotherapy. The dataset integrated multi-modal data from $n=435$ patients, comprising a total of $p=374$ variables. It included clinical and demographic variables (e.g., sex, age, $p=10$), routine blood tests (e.g., blood counts, biochemistry, $p=49$), tumor multiplex immunohistochemistry markers (e.g., immune cells, $p=159$), and circulating immune and vasculophenotyping markers. The latter included both flow cytometry data ($p=141$) and soluble ($p=15$) markers. The dataset contained 30.8% of missing values. The outcome to predict was primary resistance to immunotherapy, defined as disease progression within 6 months of treatment [39].
@@ 2.1.3 Simulated outcome
To evaluate each method’s ability to identify the "true features" (i.e., the ones hypothesized as having generated the outcome), a part of FS benchmarking implemented in roofs is conducted on a semi-synthetic dataset with a simulated outcome denoted $\tilde{y}$. With the true model being unknown, the real outcome $y$ cannot be used directly for this purpose, but the original $n\times p$ features matrix $X$ can still be used. The simulated outcome was derived by applying a logistic function to a linear combination of $S$ pre-selected features from the real dataset X with added Gaussian noise:
$\text{score}^{i}=\frac{1}{1+\exp\left(-\Big(\sum_{j\in\mathcal{S}}\beta_{j}X_{j}^{i}+\varepsilon\Big)\right)},\quad\varepsilon\sim\mathcal{N}(0,1)$
@@
where $\mathcal{S}$ is the index set of the specified true predictors and $\beta_{j}$ is the coefficient of $j^{th}$ feature estimated from an unpenalized logistic regression fitted using this set of predictors.
For the benchmark on PIONeeR data, the true predictors (19 features in total) were selected from the full dataset using a 0.01 threshold on adjusted p-values from a multivariable Cox model for progression-free survival. This approach was chosen to mimic a realistic data structure.
@@ 2.2.1 FS methods
Of the 34 FS methods implemented in roofs, 23 representative approaches were included in the present benchmark. Among embedded methods, we evaluated LASSO [44] and its variants (modified Bolasso [1] with a frequency threshold of 0.5 followed by a second LASSO fit, adaptive LASSO [50], exclusive LASSO [49], and HSIC LASSO [46]), as well as ensemble bootstrap-based methods designed to enhance stability and control false discoveries (stability selection [34], Stabl LASSO [21], RENT [26]). Both univariate and multivariate filters were used, including information theory (CIFE [33], CMIM [17], DISR [35], JMI [47]) and similarity-based approaches (ReliefF [40], Fisher score [31]), correlation-based clustering (hierarchical clustering), and statistical methods (Gini index, t-score, minimum redundancy maximum relevance (mRMR) [31], and adjusted p-values from Benjamini-Hochberg method [2]). Wrappers included two recursive elimination methods (LR-RFE, based on logistic regression coefficients, and RF-RFE, based on random forest importances), a forward selection method (forward RF), and the SHAPley-based Shapicant [7]. Finally, random feature selection was included as a control. A complete description of all benchmarked methods is provided in Supplementary Table LABEL:sup_tab1.
@@ 2.2.3 Optimism correction
$w=\frac{0.632}{1-0.368\cdot R}.$
$R$ is the overfitting ratio computed as
@@ 2.2.4 FS stability
where $s_{f}^{2}=\frac{B}{B-1}\hat{p}_{f}(1-\hat{p}_{f})$ is the unbiased sample variance of the frequency of selection $\hat{p}_{f}$ of $f^{th}$ feature over $B$ bootstraps and $\overline{k}$ is an average size of the selected subsets.
Additionally, the package reports simple measures of individual feature robustness. For each feature, it computes the selection frequency obtained by every FS method and identifies how many methods select that feature with high frequency (as defined by the user; we used a 50% threshold). The selection frequency serves as an indicator of the signal strength: features that are weakly associated with the outcome appear in only a small fraction of bootstrap models, whereas features with a strong signal are consistently selected in a large proportion of bootstrap models and across multiple FS methods. This enables to distinguish between features that are stable and predictive and those that may be artifacts of a specific algorithm or over- or undersampling of particular patient subclusters.
@@ 2.2.5 Trade-off between true and false discoveries
roofs additionally benchmarks FS methods on a semi-synthetic dataset with simulated outcome $\mathcal{D}_{\text{semi-synthetic}}=\{(x^{i},\tilde{y}^{i})\}_{i=1}^{n},\quad x^{i}\in X,\,\tilde{y}^{i}\in\tilde{Y}$. The comparison of FS methods on a semi-synthetic data allows assessing the trade-off between true and false discoveries, since the underlying true model is known. The true positive rate (TPR) – the proportion of true predictors correctly identified – represents the discovery power of an FS algorithm. The false positive rate (FPR) – the proportion of features not included in the true model but selected by an algorithm – shows the cost of discovery in terms of the error. The procedure used to derive the synthetic outcome in our benchmark on PIONeeR dataset is described in subsection 2.1.3. roofs supports modeling both linear and non-linear relationships between predictors and outcome, but in the present benchmark we used a linear model for simplicity. Variability in FPR and TPR is assessed by repeating the experiment with resampled noise, with the number of replicates specified by the user (set to 36 in our benchmark).
FS methods gathered in roofs can be subdivided into two groups based on how they determine the number of features they select: 1) methods with fixed, user-defined subset size, and 2) methods with subset size based on internal criteria, such as cross-validated lambda in the case of LASSO or p-value thresholds in the case of p.adjust or Shapicant. In the experiment on the semi-synthetic dataset, methods from the first group were configured to select 19 features, corresponding to the number of predictors in the true model. This configuration provided these methods with an advantage compared to the second group of methods with respect to both true and false positives. Their TPR was lower bounded, as by design the selection conducted by these methods was not too stringent, and only a random or very inaccurate method would have obtained a low TPR. In turn, their FPR was upper bounded, as these methods could not select more than 19 features.

### v2

@@ Abstract
Feature selection (FS) is essential for biomarker discovery and clinical predictive modeling. Over the past decades, methodological literature on FS has become rich and mature, offering a wide spectrum of algorithmic approaches. However, much of this methodological progress has not fully translated into applied biomedical research. Moreover, challenges inherent in biomedical data, such as high-dimensional feature space, low sample size, multicollinearity, and missing values, make FS non-trivial. To help bridge this gap between methodological development and practical application, we propose ROOFS (RObust biOmarker Feature Selection), a Python package available at https://gitlab.inria.fr/compo/roofs, designed to help researchers in the choice of FS method adapted to their problem. ROOFS benchmarks multiple FS methods on the user’s data and generates reports summarizing a comprehensive set of evaluation metrics, including downstream predictive performance estimated using optimism correction, stability, robustness of individual features, and true positive and false positive rates assessed on semi-synthetic data with a simulated outcome. We demonstrate the utility of ROOFS on data from the PIONeeR clinical trial, aimed at identifying predictors of resistance to anti-PD-(L)1 immunotherapy in lung cancer. Of the 34 FS methods gathered in ROOFS, we evaluated 23 in combination with 11 classifiers (253 models) and identified a filter based on the union of Benjamini-Hochberg false discovery rate-adjusted p-values from t-test and logistic regression as the optimal approach, outperforming other methods including widely used LASSO. We conclude that comprehensive benchmarking with ROOFS has the potential to improve the reproducibility of FS discoveries and increase the translational value of clinical models.
Keywords: feature selection, stability, biomarker discovery, benchmark, software
@@ 1 Introduction
Feature selection (FS) refers to a class of methods aimed at reducing the feature space by selecting a subset of size $k<p$, where $p$ denotes the number of features. In supervised learning, FS is used either for discovery, to identify the most informative variables, or for prediction, to reduce variance. Over the past decades, a wide range of FS methods has been proposed (see [20] and [34] for extensive reviews). FS methods can be classified into 4 groups: filter, embedded, wrapper, and ensemble methods. Filters evaluate features according to a specific criteria and output either a complete ranking or a subset selected using a predefined threshold (e.g., a statistical test). Embedded methods perform FS simultaneously with model learning, as in regularization-based approaches or tree-based models with constrained depth. Wrapper methods iteratively search for feature subsets that maximize the predictive performance of a given model. Ensemble methods aggregate outputs from different selection algorithms or from repeated runs of a single method on resampled data. Filter and embedded methods are typically more straightforward, computationally efficient, and interpretable compared to wrapper and ensemble techniques. However, computationally expensive and algorithmically complex wrapper methods possess an important property: an ability to capture feature interactions tailored to a specific classifier, typically resulting in modest performance improvement. The computational cost of the ensemble methods is balanced by their ability to select features with the strongest signal and avoid false discoveries.
A basic requirement for FS is that the reduced model should perform equivalently to or better than the full model, capturing the complexity of the original data. In the context of biomarker discovery and clinical predictive models, the selected subset, often called a “signature“, should meet additional requirements: (i) contain biologically or clinically interpretable features; (ii) be as minimal as possible; and (iii) demonstrate stability, i.e. the selected features should not be sensitive to small changes in the data [31]. FS stability is of key importance for ensuring reproducibility. It is typically assessed by measuring the similarity between feature subsets obtained across resamples of the original data, e.g. via bootstrapping or subsampling. A method that is not stable may result in models that do not generalize well to external populations, unlikely to be clinically useful.
Feature selection is a long-standing field of research in machine learning [20]. However, much of the methodological progress has not fully translated into applied biomedical research, where practitioners often rely on at most a limited set of familiar methods. Likewise, bootstrap-based procedures for assessing stability are rarely used, despite its importance having been frequently acknowledged [31, 23, 25]. In addition, the systematic evaluation of the suitability of FS methods for a given dataset is often omitted, although previous benchmark studies [5, 15] have shown that, in terms of predictive performances, the optimal method is contingent upon the specific dataset. Furthermore, FS performance is sensitive to hyperparameter choices, such as the selected subset size, and different settings can lead to different rankings of methods on the same dataset [48].
@@
To address this gap between methodological and applied research, we propose a Python package called ROOFS (RObust biOmarker Feature Selection), available at https://gitlab.inria.fr/compo/roofs. It conducts comprehensive benchmarking of multiple pre-implemented FS algorithms on the user’s data based on three criteria: 1) method stability and robustness of individual features, 2) predictive performances, and 3) true positive and false positive rates (from semi-synthetic data with a simulated outcome). For point 2), we relied on the optimism-correction (OC) framework emphasized as a guideline for the development of clinical predictive models by Collins et al. [14]. ROOFS provides automated reporting of the evaluation metrics listed above – a key feature of the package aimed to facilitate comparison and choice of an FS method adapted to the user’s problem.
@@
We demonstrate the practical value of ROOFS by applying it to a complex dataset from the PIONeeR clinical study (NCT03833440, Precision Immuno-Oncology for Advanced Non-small Cell Lung Cancer Patients With PD-1 ICI Resistance) to predict primary resistance to immunotherapy. This dataset is an interesting and challenging use case for FS because: 1) the data is rich and heterogeneous combining clinical features, routine blood tests and detailed immunoprofiling of blood and tumor samples measured by different clinical and research partners; 2) it contains clusters of highly correlated features, e.g. numerous variables describing related immune cell populations; and 3) $p$ (374) is roughly equivalent to $n$ (435). This is different from the three types of datasets popular in the literature: a) high-dimensional omics ($p\gg n$); b) data with a sufficiently high event-per-variable ratio, e.g. the number of patients in the smallest class is at least ten times the number of predictors; and c) big data (e.g., from electronic health records), with $n\gg p$ typically modeled using deep artificial neural networks.
@@
A comprehensive survey of FS benchmarks published before 2020 across multiple domains is provided by Hopf and Reifenrath [26]. Table 1 summarizes more recent benchmarks on high-dimensional biomedical data. Most of these studies focused on a single data modality (genomic data). Only two studies included FS methods representing all algorithmic families, and stability was evaluated in only three of the eight benchmarks. The novelty of our study are that: 1) it addresses heterogeneous, mixed-type data; 2) it evaluates a broader and more diverse range of FS algorithms; 3) it reports all standard performance metrics, with an emphasis on stability; and 4) it extends FS evaluation to assess the balance between true and false discoveries.
@@ F E W Ens Predictive performance Stability Runtime
Our work Cancer multi-type 1 23
@@ 2.1.1 PIONeeR dataset
The development of ROOFS was motivated by the following real-world problem: identifying robust biomarkers of resistance to immunotherapy in lung cancer. To this end, we relied on the PIONeeR clinical trial dataset, derived from a prospective, multicenter cohort of patients (17 centers in France) with advanced or recurrent non-small cell lung cancer (NSCLC) [2]. Patients received either frontline combination therapy with platinum-based chemotherapy plus anti-PD-(L)1 immune checkpoint inhibitors (ICIs) or second-/third-line monotherapy with anti-PD-(L)1 ICIs following progression on prior platinum-based chemotherapy. The dataset integrated multi-modal data from $n=435$ patients, comprising a total of $p=374$ variables. It included clinical and demographic variables (e.g., sex, age, $p=10$), routine blood tests (e.g., blood counts, biochemistry, $p=49$), tumor multiplex immunohistochemistry markers (e.g., immune cells, $p=159$), and circulating immune and vasculophenotyping markers. The latter included both flow cytometry data ($p=141$) and soluble ($p=15$) markers. The dataset contained 30.8% of missing values. The outcome to predict was primary resistance to immunotherapy, defined as disease progression within 6 months of treatment [45].
@@ 2.1.3 Simulated outcome
To evaluate each method’s ability to identify the "true features" (i.e., the ones hypothesized as having generated the outcome), a part of FS benchmarking implemented in ROOFS is conducted on a semi-synthetic dataset with a simulated outcome denoted $\tilde{y}$. With the true model being unknown, the real outcome $y$ cannot be used directly for this purpose, but the original $n\times p$ features matrix X can still be used. The synthetic outcome $\tilde{y}^{(i)}$ for observation $i$ is generated by applying a logistic function to a linear combination of $S$ pre-selected features from the real dataset X with added Gaussian noise:
$\text{score}^{i}=\frac{1}{1+\exp\left(-\Big(\sum_{j\in\mathcal{S}}\beta_{j}X_{j}^{i}+\varepsilon^{i}\Big)\right)},\quad\varepsilon^{i}\sim\mathcal{N}(0,1)$
@@
where $\mathcal{S}$ is the index set of the pre-selected true predictors and $\beta_{j}$ is the coefficient of the $j^{th}$ predictor. The coefficients are determined using one of the two approaches: 1) Random sampling (default approach in ROOFS): sample from $N(\mu,\sigma^{2})$ and randomly assign the sign (the default parameters $\mu=2$ and $\sigma=0.5$); 2) Estimation from the data: fit an unpenalized logistic regression on the true outcome $y$ with this set of predictors and set $\beta_{j}$ equal to the estimated coefficient for predictor $j$.
For the benchmark on PIONeeR data, 19 true predictors were selected from the full dataset using a 0.01 threshold on Benjamini-Hochberg adjusted p-values from a multivariable Cox model for progression-free survival. This approach was chosen to mimic a realistic data structure. To preserve the relative effect magnitudes of the true predictors, we used approach (2) above to derive the coefficients.
The approach outlined above uses a linear model for simplicity, but ROOFS also supports nonlinearity through polynomial power terms and pairwise feature interactions.
@@ 2.2.1 FS methods
Of the 34 FS methods implemented in ROOFS, 23 representative approaches were included in the present benchmark. Among embedded methods, we evaluated LASSO [49] and its variants (adaptive LASSO [55] and exclusive LASSO [54]), as well as ensemble bootstrap-based methods designed to enhance stability and control false discoveries (modified Bolasso [1] with a frequency threshold of 0.5 followed by a second LASSO fit, stability selection [37], Stabl LASSO [24], RENT [29], and HSIC LASSO [51]). Both univariate and multivariate filters were used, including information theory-based techniques (CIFE [36], CMIM [19], DISR [38], JMI [52]) and similarity-based approaches (ReliefF [46], Fisher score [34]), correlation-based clustering (hierarchical clustering), and statistical methods (Gini index, t-score, minimum redundancy maximum relevance (mRMR) [34], and adjusted p-values using Benjamini-Hochberg method [3]). Wrappers included two recursive elimination methods (LR-RFE, based on logistic regression coefficients, and RF-RFE, based on random forest importances), a forward selection method (forward RF), and the SHAPley-based Shapicant [8]. Finally, random feature selection was included as a control. A complete description of all benchmarked methods is provided in Supplementary Table LABEL:sup_tab1.
@@ 2.2.3 Optimism correction
$w=\frac{0.632}{1-0.368\cdot R}$
and $R$ is the overfitting ratio computed as:
@@ 2.2.4 FS stability
where $s_{f}^{2}=\frac{B}{B-1}\hat{p}_{f}(1-\hat{p}_{f})$ is the unbiased sample variance of the frequency of selection $\hat{p}_{f}$ of $f^{th}$ feature over $B$ bootstraps and $\overline{k}$ is an average size of the selected subsets. $S$ takes values from $\frac{1}{B-1}$ to 1, with the lower bound asymptotically approaching 0 as $B$ approaches infinity. $S=1$ indicates perfect stability when the same features are selected at each iteration. $S=0$ corresponds to the case of random selection.
Additionally, ROOFS reports simple measures of individual feature robustness. For each feature, it computes the selection frequency obtained by every FS method and identifies how many methods select that feature with high frequency (as defined by the user; we used a 50% threshold). The selection frequency serves as an indicator of the signal strength: features that are weakly associated with the outcome appear in only a small fraction of bootstrap models, whereas features with a strong signal are consistently selected in a large proportion of bootstrap models and across multiple FS methods. This enables to distinguish between features that are stable and predictive and those that may be artifacts of a specific algorithm or over- or undersampling of particular patient subclusters.
@@ 2.2.5 Trade-off between true and false discoveries
ROOFS additionally benchmarks FS methods on a semi-synthetic dataset with simulated outcome $\mathcal{D}_{\text{semi-synthetic}}=\{(x^{i},\tilde{y}^{i})\}_{i=1}^{n},\quad x^{i}\in X,\,\tilde{y}^{i}\in\tilde{Y}$. The comparison of FS methods on a semi-synthetic data allows assessing the trade-off between true and false discoveries, since the underlying true model is known. The true positive rate (TPR) – the proportion of true predictors correctly identified – represents the discovery power of an FS algorithm. The false positive rate (FPR) – the proportion of features not included in the true model but selected by an algorithm – shows the cost of discovery in terms of error. The procedure used to derive the synthetic outcome in our benchmark on PIONeeR dataset is described in subsection 2.1.3. The variability in FPR and TPR is assessed by repeating the experiment with resampled noise, with the number of replicates specified by the user (set to 40 in our benchmark).
FS methods gathered in ROOFS can be subdivided into two groups based on how they determine the number of features they select: 1) methods with fixed, user-defined subset size, and 2) methods with varying subset size dependent on internal criteria, such as cross-validated lambda in the case of LASSO or p-value thresholds in the case of p.adjust or Shapicant. In the experiment on the semi-synthetic dataset, methods from the first group were configured to select 19 features, corresponding to the number of predictors in the true model. This configuration provided these methods with an advantage compared to the second group of methods with respect to both true and false positives. Their TPR was lower bounded, as by design the selection conducted by these methods was not too stringent, and only a random or very inaccurate method would have obtained a low TPR. In turn, their FPR was upper bounded, as these methods could not select more than 19 features.